### v1

# The Initial Value Problem for Harmonic maps of Cohomogeneity One manifolds

Abstract.
We set up and solve the initial value problem for equivariant harmonic maps of cohomogeneity one manifolds, i.e. we show the local existence of a harmonic map in the neighborhood of a singular orbit. Furthermore, we present some theory of regular-singular systems of first order.
Key words and phrases:
harmonic map; initial value problem
2010 Mathematics Subject Classification:
58E20; 53C43
Anna Siffert thanks the German Research Foundation (DFG) for funding - Project-ID 427320536 - SFB 1442. This project is part of Anna Siffert’s project on the construction of harmonic maps between cohomogeneity one manifolds from the SFB-funding period 2020-2024.
1. Introduction
The study of harmonic maps between Riemannian manifolds and their generalizations is an active research area within geometric analysis, see e.g. the recent manuscripts [5, 6] and the references therein as well as the book [3].
A smooth map $\varphi:(M,g)\rightarrow(N,h)$ between Riemannian manifolds is called harmonic if it is a critical point of the energy functional. In what follows we explain the definition of harmonic maps with more details, where we closely follow [12]. Let $\Omega\subset M$ be a domain with piecewise $C^{1}$-boundary $\partial\Omega$. The energy integral of $\varphi$ over the domain $\Omega$ is defined as
$E_{\Omega}(\varphi)=\frac{1}{2}\int_{\Omega}\lvert d\varphi\rvert^{2}v_{g},$ (1.1)
where $v_{g}$ denotes the volume form of $M$ with respect to the metric $g$. Let $\Phi:M\times(-\varepsilon,\varepsilon)\rightarrow N,(x,t)\mapsto\Phi(x,t)=:\varphi_{t}(x)$, $\varepsilon\in\mathbb{R}$ with $\varepsilon>0$, be a smooth map with $\Phi(x,0)=\varphi(x)$, i.e. $\Phi$ is a one-parameter variation of $\varphi$. The map $\Phi$ is called supported in $\Omega$, if $\varphi_{t}=\varphi$ on $M\setminus\Omega^{\circ}$, where $\Omega^{\circ}$ denotes the interior of $\Omega$. A smooth map $\varphi:(M,g)\rightarrow(N,h)$ is called harmonic map if it is critical point of (1.1), i.e. for all compact domains $\Omega$ and all smooth one-parameter variations $\Phi$ of $\varphi$ supported in $\Omega$, we have
$\frac{d}{dt}E_{\Omega}(\varphi_{t})_{\lvert t=0}=0,$
i.e. the first variation of the energy functional vanishes.
Straightforward computations yield that the first variation of the energy functional is given by
$\frac{d}{dt}E_{\Omega}(\varphi_{t})_{\lvert t=0}=-\int_{M}\langle\tau(\varphi),\frac{\partial\varphi_{t}}{\partial t}_{\lvert t=0}\rangle v_{g},$
where $\langle\,\cdot\,$$\cdot\,\rangle$ denotes the inner product on $\varphi^{-1}TN$ which is induced from $h$ and the tension field of $\varphi$, denoted by $\tau(\varphi)$, is defined by
$\tau(\varphi)=\sum_{i=1}^{\dim M}({}^{\varphi}\nabla_{e_{i}}(d\varphi\,e_{i})-d\varphi(\nabla_{e_{i}}e_{i})).$
Here $\nabla$ denotes the Levi-Civita connection of $(M,g)$, $^{\varphi}\nabla$ is the pull-back of the Levi-Civita connection of $(N,h)$ to the bundle $\varphi^{-1}TN$ and $(e_{i})_{i=1}^{\dim M}$ is a local orthonormal frame on $M$. Clearly, $\varphi$ is harmonic if and only if
$\tau(\varphi)=0.$ (1.2)
In local coordinates the identity (1.2) constitutes a semilinear elliptic system of partial differential equations of second order. There is no general solution theory for such systems, providing solutions to the equation (1.2) is typically a hard task. A common strategy to construct solutions is to impose symmetry conditions such that the original problem reduces to a simpler, in best cases to a (partially) solvable, problem. One difficulty, of course, is finding the right symmetry conditions to generate a problem that is (partially) solvable.
One fruitful symmetry assumption in the context of harmonic maps is to study equivariant maps between cohomogeneity one manifolds for harmonicity. The study of equivariant harmonic maps between cohomogeneity one manifolds was initiated by Urakawa [22] and has been widely extended by Püttmann and Siffert in [17]. Using this Ansatz, examples of harmonic maps between cohomogeneity one manifolds have been provided for specific choices of these cohomogeneity one manifolds, see e.g. [4, 17, 19, 20, 22].
Recall that $M$ is called a cohomogeneity one manifold if it supports a smooth isometric action $G\times M\rightarrow M$ of a compact Lie group $G$ such that the orbit space $M/G$ has dimension one. There are two types of orbits: Regular and non-regular orbits. Regular orbits have codimension one and their union constitutes an open dense set $M_{0}$ in $M$. A non-regular orbit with codimension strictly greater than one is called singular orbit. Boundary points of $M/G$ correspond to non-regular orbits, more precisely the non-regular orbits are the fibres of the boundary points of $M/G$ under the natural projection $\pi:M\rightarrow M/G$. Recall further that equivariant maps are structure preserving maps between cohomogeneity one manifolds, i.e. they map orbits to orbits.
For equivariant maps $\varphi$ between cohomogeneity one manifolds equation (1.2) reduces to a singular boundary value problem. More precisely one obtains an ordinary differential equation of second order for a smooth function $r:(M/G)^{\circ}\rightarrow\mathbb{R}$ of the form
$\ddot{r}(t)+f_{1}(t)\dot{r}(t)+f_{2}(t,r(t))=0,$ (1.3)
where $h_{1}:(M/G)^{\circ}\rightarrow\mathbb{R},\,h_{2}:(M/G)^{\circ}\times\mathbb{R}\rightarrow\mathbb{R}$ are smooth maps and $(M/G)^{\circ}$ denotes the interior of $M/G$. In order for $\varphi$ to be harmonic, $r$ has to satisfy specific boundary conditions at the boundary points of $M/G$. At boundary points of $M/G$ corresponding to singular orbits the differential equation (1.3) becomes singular. At points of $M/G$ whose fibres under the natural projection $\pi:M\rightarrow M/G$ are not singular orbits, the differential equation (1.3) is regular. To summarize, in order to construct equivariant harmonic maps between cohomogeneity one manifolds one has to provide solutions of (1.3) which extend smoothly to the boundary points of $M/G$. In this manuscript we will solve the associated initial value problem, i.e. we assume that $0\in M/G$ is a boundary point and show that for each $v$ there exists a unique smooth solution $r:(M/G)^{\circ}\cup\{0\}\rightarrow\mathbb{R}$ of (1.3) with $r(0)=0$ and $\dot{r}(0)=v$ which depends continuously on $v$. Since this is a local problem it is sufficient to study this problem near one singular orbit only.
Theorem 1.1.
Let a cohomogeneity one manifold $M$ with a singular orbit $N$ be given. Assume that $N$ is the fibre of $0\in M/G$ under the natural projection map $\pi:M\rightarrow M/G$. For each $v$ there exists a unique smooth solution $r:(M/G)^{\circ}\cup\{0\}\rightarrow\mathbb{R}$ of (1.3) with $r(0)=0$ and $\dot{r}(0)=v$ which depends continuously on $v$.
A precise formulation of this result is contained in Section 3. For specific cases the initial value problem for harmonic maps of cohomogeneity one manifolds has been solved in the literature, see e.g. [7, 10, 19, 20].
In Section 4 we study regular-singular systems of first order. Such systems appear naturally in the context of the initial value problem for harmonic maps between cohomogeneity one manifolds. A lot of the presented results in Section 4 are known, but they are somehow scattered in the literature. Here we provide an overview that will hopefully also be useful for future research works.
Organization: In Section 2 we provide preliminaries. The initial value problem for harmonic maps of cohomogeneity one manifolds is set up and solved in Section 3. In Section 4 we study regular-singular systems of first order.
2. Preliminaries
We provide preliminaries on cohomogeneity one manifolds, on smoothness conditions for cohomogeneity one metrics and on harmonic maps between Riemannian manifolds.
2.1. Cohomogeneity one manifolds
In this subsection we briefly recall cohomogeneity one manifolds. This subsection closely follows the exposition in [17], a few sentences have been copied almost word for word.
Throughout let $M$ be a connected manifold. $M$ has cohomogeneity one if it supports a smooth isometric action $G\times M\rightarrow M$ of a compact Lie group $G$ such that the orbit space $M/G$ has dimension one. In this case, $M$ is also referred to simply as cohomogeneity one manifold.
Mostert [15, 16] proved a classification result of cohomogeneity one manifolds, namely he showed that the orbit space $M/G$ of such manifolds is isometric to exactly one of the following spaces:
  1. (1)
    the real line, i.e. $M/G=\mathbb{R}$;
  2. (2)
    the circle, i.e. $M/G=S^{1}$;
  3. (3)
    a closed interval, i.e. $M/G=[0,L]$ for some $L\in\mathbb{R}_{+}$;
  4. (4)
    a half line, i.e. $M/G=[0,\infty)$.
We denote by $\partial(M/G)$ the boundary points of the orbit space $M/G$, i.e. $\partial(\mathbb{R})=\emptyset$, $\partial(S^{1})=\emptyset$, $\partial([0,L])=\{0,L\}$ and $\partial([0,\infty))=\{0\}$.
Below let $M$ be a cohomogeneity one manifold. Each boundary point, i.e. each element of $\partial(M/G)$, corresponds to a non-principal orbit $N$. In case (3) there are exactly two non-principal orbits, which we denote by $N_{0}$ and $N_{1}$, respectively. In case (4) there is exactly one non-principal orbit, which we denote by $N_{0}$. Further, each interior point of $M/G$ corresponds to a principal orbit.
A non–principal orbit $N$ is called exceptional orbit, or simply exceptional, if its dimension equals the common dimension of the principal orbits. Otherwise we call $N$ singular orbit, or simply singular. The collection $M_{0}$ of all principal orbits is called the regular part of $M$.
Let $\gamma$ be a unit-speed normal geodesic, this means a geodesic $\gamma:\mathbb{R}\to M$ that passes through all orbits perpendicularly and satisfies $\gamma(0)\in N_{0}$ and $\gamma(L)\in N_{1}$ in case (3), and $\gamma(0)\in N_{0}$ in case (4). The isotropy groups of the regular points $\gamma(t)$, $t\in(M/G)^{\circ}$, are constant. This common principal isotropy group is henceforth denoted by $H$. The point of introducing normal geodesics of cohomogeneity one manifolds is that geometric problems on cohomogeneity one manifolds can be reduced to the study of these geometric problems along an arbitrary but fixed normal geodesic.
In case (3) we further assume that $\gamma$ is closed or, equivalently, that the Weyl group $W$ is finite. Recall that the Weyl group $W$ is defined to be the subgroup of the elements of $G$ that leave $\gamma$ invariant modulo the subgroup of elements that fix $\gamma$ pointwise. It is a dihedral subgroup of $N(H)/H$ generated by two involutions that fix $\gamma(0)$ and $\gamma(L)$, respectively, and it acts simply transitively on the regular segments of the normal geodesic.
In this manuscript we study the local structure of a cohomogeneity one manifold near a singular orbit $N$. Let $\gamma$ be a unit speed normal geodesic such that $x_{0}:=\gamma(0)\in N$. Denote the isotropy group at $x_{0}$ by $K$. Thus $N=G\cdot x_{0}=G/K$.
The Riemannian metric $g$ of $M$ is described in terms of a one parameter family of $G$-invariant metrics $g_{t}$ on the principal orbits $G/H$, i.e. we have
$g=g_{t}+dt^{2},$ (2.1)
where $g_{t}$ has to satisfy invariance conditions in case (2) and smoothness conditions in cases (3) and (4), see e.g. [11] and the references therein. The smoothness conditions for cases (3) and (4) are discussed in the following subsection. Note that since $M_{0}$ is dense in $M$, (2.1) also describes the metric on the whole of $M$.
2.2. Smoothness conditions of cohomogeneity one metrics
Suppose that $M$ is a cohomogeneity one manifold with (at least one) singular orbit. Assume further that the dense and open set $M_{0}\subset M$ is endowed with a $G$-invariant metric. In [8] and [24] it has been investigated under which conditions the extension of this metric to the singular orbit is smooth and smoothness conditions for cohomogeneity one metrics have been provided. The description of smoothness conditions given by Verdiani and Ziller in [24] is often easier to apply in practice, compare [25]. However, for our purposes the results of [8] are sufficient. The following presentation closely follows [24].
While a non-compact cohomogeneity one manifold with a singular orbit is a vector bundle, such a compact cohomogeneity one manifold consists of the union of two vector bundles, see e.g. [1, 2] and the references therein. In this manuscript we will study the local structure of a cohomogeneity one manifold near a singular orbit. We can thus restrict ourselves to one vector bundle.
Let $G$ be a compact Lie group and $H,G\subset G$ be closed Lie subgroups such that $H\subset K\subset G$ and $K/H=\mathbb{S}^{\ell}$ for a $\ell\in\mathbb{N}=\{1,2,3,\dots\}$. In other words, $K$ acts transitively on $\mathbb{S}^{\ell}$, i.e. there exists a representation $\rho:K\rightarrow O(\ell+1)$ such that $\rho(K)x=\mathbb{S}^{\ell}$ for each $x\in\mathbb{S}^{\ell}$. The action of the isotropy group $K$ on the sphere $\mathbb{S}^{\ell}$ extends to a linear action on the disk $D:=D^{\ell+1}\subset\mathbb{R}^{\ell+1}$. Therefore $M=G\times_{K}D$ is a homogeneous disc bundle, with base space $G/K$ and structure group $H$. Here the isotropy group $K$ acts on $G\times D$ as follows
$(k,(g,d))\mapsto(gk^{-1},\rho(k)d).$
Note further that $G$ acts with cohomogeneity one on $M$ by left multiplication, i.e.
$(g,[\hat{g},d])\mapsto[g\hat{g},d]$
for $d\in D$ and $g,\hat{g}\in G$. The singular orbit $N:=G\times_{K}\{0\}$ is diffeomorphic to $G/K$. The pricincipal orbits are given by $P_{t}:=G\times_{K}\mathbb{S}^{\ell}_{t}$ for $t$ small enough. Here $\mathbb{S}^{\ell}_{t}\subset D$ denotes the sphere of radius $t$. Each $P_{t}$ is diffeomorphic to $G/H$. Using the exponential map, the disk $D$ can G equivariantly be identified with a manifold which is orthogonal to the singular orbit at the point $\gamma(0)$. Thus $D$ can be considered as slice of the $G$ action on $M$.
To state the main result of [24], let $\gamma\colon[0,\infty)\to\mathbb{R}^{\ell+1}$ be a normal geodesic. The metric on $M_{0}$ is determined by its values along $\gamma$. This follows since these values and the action of $G$ determine the metric on the whole of $M_{0}$. Let $\mathfrak{g},\mathfrak{h}$ be the Lie algebras of $G$ and $H$, respectively. Let $Q$ be a fixed biinvariant metric on $G$. Further, let $\mathfrak{n}$ be the orthonormal complement of the Lie algebra $\mathfrak{h}$ in $\mathfrak{g}$. The complement $\mathfrak{n}$ can be identified with the tangent space to the regular orbits along $\gamma$ by action fields, i.e. $X\in\mathfrak{n}\mapsto X^{\ast}(\gamma(t))$.
Let $X_{i}$ be a basis of $\mathfrak{n}$. Then $X^{*}_{i}(\gamma(t))$ is a basis of $\dot{\gamma}^{\perp}(t)\subset T_{\gamma(t)}M$ for each $t>0$. Further, the metric is determined by the collection of the $r$ functions $g_{ij}(t)=g(X_{i}^{*},X_{j}^{*})_{c(t)},\ i\leq j$.
Wit this preparation at hand we can cite the following theorem which provides smoothness conditions for $g$:
Theorem 2.1 ([24]).
Let $g_{ij}(t),\ t>0$ be a smooth family of positive definite matrices describing the cohomogeneity one metric on the regular part along a normal geodesic $\gamma(t)$. Then there exist integers $a_{ij}^{k}$ and $d_{k}$, with $d_{k}\geq 0$, such that the metric has a smooth extension to all of $M$ if and only if
$\sum_{i,j}a_{ij}^{k}\,g_{ij}(t)=t^{d_{k}}\phi_{k}(t^{2})\quad\text{ for }k=1,\cdots,r,\ \text{ and }t>0$
where $\phi_{1},\cdots,\phi_{r}$ are smooth functions defined for $t\geq 0$.
Let $Q$ be a fixed biinvariant metric on $G$. With respect to $Q$ we consider the following splitting of the Lie algebra $\mathfrak{g}$:
$\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}\oplus\mathfrak{m}.$
Here $\mathfrak{h}$ is the Lie algebra of $H$ and $\mathfrak{h}\oplus\mathfrak{p}$ is the Lie algebra of $K$.
Theorem 2.1 in particular yields
$g_{\lvert\mathfrak{m}}=A_{0}+tA_{1}+t^{2}A(t),$
$g_{\lvert\mathfrak{p}}=t^{2}\operatorname{Id}_{\mathfrak{p}}+t^{4}B(t),$
$g_{\lvert\mathfrak{pm}}=t^{2}C_{0}+t^{3}C(t),$
where $A_{0}$, $A_{1}$ and $C_{0}$ are constant matrices and $A,B,C$ are smooth matrix valued functions of $t$, see [25].
2.3. Harmonic maps between cohomogeneity one manifolds
We study equivariant harmonic maps between cohomogeneity one manifolds. Again we borrow notation and inspiration for presentation from [17]. For simplicity of the presentation we restrict ourselves to self-maps. Note however that the results of this subsection can be easily extended to maps between possibly different cohomogeneity one manifolds.
Equivariant maps are the structure preserving maps between cohomogeneity one manifolds, i.e. these maps map orbits to orbits. In case (3), Püttmann [18] constructed an infinite family of equivariant smooth self-maps of $M$ given by
$\psi:M\rightarrow M,\,g\cdot\gamma(t)\mapsto g\cdot\gamma(kt).$
Here $\gamma$ is a unit speed normal geodesic such that $\gamma(0)\in N_{0}$. Furthermore, $k\in\mathbb{Z}$ is of the form $j|W|/2+1$ where $j\in 2\mathbb{Z}$. Under specific conditions on the action $G\times M\rightarrow M$ odd integers $j$ might also be allowed, see Lemma 2.1 in [18]. Following [17], we the map $g\cdot\gamma(t)\mapsto g\cdot\gamma(kt)$ is called the $k$-map of $M$. If the codimensions of the non-principal orbits $N_{0}$ and $N_{1}$ are both odd the degree of a $k$-map is equal to $k$. If the codimensions of the non-principal orbits $N_{0}$ and $N_{1}$ are not both odd degree of a $k$-map is equal to 0 or $\pm 1$.
Below let $M$ be an arbitrary cohomogeneity one manifold, i.e. we do not assume any longer that we are necessarily dealing with case (3), and let $\gamma$ be a fixed but arbitrary normal geodesic.
Below we study the equivariant self-maps of $M$ given by
$\psi:M\rightarrow M,\,g\cdot\gamma(t)\mapsto g\cdot\gamma(r(t)).$ (2.2)
Here $r:M/G\rightarrow\mathbb{R}$ is a smooth function which in addition satisfies boundary conditions in cases (3) and (4). Namely, $r$ satisfies
  1. (i)
    the boundary conditions $r(0)=0$ and $r(L)=kL$, in case $(3)$. Here $k=j|W|/2+1$ where $j\in 2\mathbb{Z}$ (and possibly $j\in\mathbb{Z}$, see above);
  2. (ii)
    the boundary condition $r(0)=0$ in case $(4)$.
Analogous to Lemma 2.1 in [18], one proves that in case (4) the maps (2.2) are smooth. Thus, the maps (2.2) are smooth in all cases (1)-(4).
In case (3), Püttmann and Siffert [17] studied the maps (2.2) for harmonicity. For this purpose the tension field of $\psi$ was calculated in case $(3)$, see Theorems A and B in [17]. The calculations in the proofs of Theorems A and B are local. We can thus conduct the same calculations in cases (1), (2) and (3).
In order to state these results, we introduce the following notation, which was also used in [17]. By
$\Pi_{t}^{r(t)}:T_{\gamma(t)}(G\cdot\gamma(t))\to T_{\gamma(r(t))}(G\cdot\gamma(r(t))$
we denote the parallel transport along the normal geodesic $\gamma$. The action field homomorphism, i.e. the map given by $X^{\ast}_{|\gamma(t)}\mapsto X^{\ast}_{|\gamma(r(t))}$, constitutes a homomorphism between the two tangent spaces $T_{\gamma(t)}(G\cdot\gamma(t))$ and $T_{\gamma(r(t))}(G\cdot\gamma(r(t)))$. The endomorphism of $T_{\gamma(t)}(G\cdot\gamma(t))$ given by composing the action field homomorphism with $(\Pi_{t}^{r(t)})^{-1}=\Pi_{r(t)}^{t}$ is denoted by $J_{t}^{r(t)}$.
We get the following result for the normal component of the tension field of the maps $\psi$ given in (2.2) (compare Theorem A in [17]):
Theorem 2.2.
The normal component of the tension field of $\psi$, see (2.2), is given by
$\tau^{\mathrm{nor}}_{|\gamma(t)}=\ddot{r}(t)-\dot{r}(t)\operatorname{Tr}S_{|\gamma(t)}+\operatorname{Tr}\,(J_{t}^{r(t)})^{\ast}(\Pi_{t}^{r(t)})^{-1}S_{|\gamma(r(t))}\Pi_{t}^{r(t)}J_{t}^{r(t)}$
for $t\in(M/G)^{\circ}$. Here, $S_{|\gamma(t)}$ denotes the shape operator of the orbit $G\cdot\gamma(t)$ at $\gamma(t)$ and $(J_{t}^{r(t)})^{\ast}$ denotes the adjoint endomorphism of $J_{t}^{r(t)}$.
Recall that in the previous subsection we introduced $Q$, a fixed biinvariant metric on $G$. We denote by $\mathfrak{n}$ the orthonormal complement of the Lie algebra $\mathfrak{h}$ of the principal isotropy group $H$ in $\mathfrak{g}$. We define the metric endomorphisms $P_{t}:\mathfrak{n}\to\mathfrak{n}$ by
$Q(X,P_{t}\cdot Y)=g(X^{\ast},Y^{\ast})_{\gamma(t)}.$
With this notation the normal component of the tension field can be expressed as follows:
Theorem 2.3 (Theorem 3.4 in [17]).
The normal component of the tension field is given by
$\tau^{\mathrm{nor}}_{|\gamma(t)}=\ddot{r}(t)+\tfrac{1}{2}\dot{r}(t)\operatorname{Tr}P_{t}^{-1}\dot{P}_{t}-\tfrac{1}{2}\operatorname{Tr}P_{t}^{-1}(\dot{P})_{r(t)}.$
We obtain the following result for the tangential component of the tension field of $\psi$ (compare Theorem B in [17]):
Theorem 2.4.
The tangential component of the tension field of a $\psi$ is given by
$\tau^{\tan}_{|\gamma(t)}=-\sum_{\mu,\nu=1}^{n}\langle[E_{\mu},F_{\nu}]^{\ast},E_{\mu}^{\ast}\rangle_{|\gamma(r(t))}F^{\ast}_{\nu|\gamma(r(t))}$
for $t\in(M/G)^{\circ}$. Here, $E_{1},\ldots,E_{n}\in\mathfrak{n}$ and $F_{1},\ldots,F_{n}\in\mathfrak{n}$ are such that $E^{\ast}_{1|\gamma(t)}$,…, $E^{\ast}_{n|\gamma(t)}$ form an orthonormal basis of $T_{\gamma(t)}(G\cdot\gamma(t))$ and $F^{\ast}_{1|\gamma(r(t))},\ldots,F^{\ast}_{n|\gamma(r(t))}$ form an orthonormal basis of $T_{\gamma(r(t))}(G\cdot\gamma(r(t)))$.
3. The initial value problem for harmonic maps
In this section we set up and solve the initial value problem for harmonic maps of cohomogeneity one manifolds. We will assume that a singular orbit $N$ exists, otherwise the initial value problem is a regular ODE. Thus $M/G$ is either isometric to $[0,1]$ or $0,\infty)$. Below we set up the initial value problem at $t=0$.
In this section we solve the initial value problem for harmonic maps at singular orbits:
Theorem 3.1.
Let $(M,g)$ be a cohomogeneity one manifold with singular orbit, i.e. $M/G$ is isometric to a closed interval $[0,L]$ or a half line $0,\infty)$. For any $v\in\mathbb{R}$ the differential equation
$0=\ddot{r}(t)-\dot{r}(t)\operatorname{Tr}S_{|\gamma(t)}+\operatorname{Tr}\,(J_{t}^{r(t)})^{\ast}(\Pi_{t}^{r(t)})^{-1}S_{|\gamma(r(t))}\Pi_{t}^{r(t)}J_{t}^{r(t)}$
has a smooth solution $r:(M/G)^{\circ}\cup\{0\}\to\mathbb{R}$ with $\lim_{t\to 0}r(t)=0$ and $\lim_{t\to 0}\dot{r}(t)=v$, which depends continously on $v$.
Before proving Theorem 3.1 we recall the following: We consider metrics of the form (2.1) and have set
$g(X^{\ast},Y^{\ast})_{\gamma(t)}=g_{t}(X,Y)=Q(P_{t}X,Y)$
for $X,Y\in\mathfrak{n}$. Note that the endomorphism $P_{t}:\mathfrak{n}\rightarrow\mathfrak{n}$ is $\operatorname{Ad}_{H}$-equivariant and $Q$-symmetric. With respect to the metric $Q$ we consider the following splitting of the Lie algebra $\mathfrak{g}$:
$\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}\oplus\mathfrak{m}.$
Here $\mathfrak{h}\oplus\mathfrak{p}$ is the Lie algebra of $K$. Note further that
$\mathfrak{n}=\mathfrak{p}\oplus\mathfrak{m}.$
Note that $H$ acts on $\mathfrak{n}$ via the adjoint representation. Further, a $G$ invariant metric on $G/H$ is described by an $\operatorname{Ad}_{H}$ invariant inner product on $\mathfrak{n}$. [^1] The notation and facts summarized in the previous paragraph have been used in various manuscripts, see e.g. [11, 25, 24, 17].
We are now ready to prove Theorem 3.1.
Proof of Theorem 3.1.
The first step of the proof consists in determining the initial value problem.
As in the manuscripts [25, 8] we first make the following assumption:
Assumption 3.2.
There are no $\operatorname{Ad}_{H}$ irreducible representations in $\mathfrak{m}$ which are equivalent to one in $\mathfrak{p}$.
We will remove this assumption later in this proof. It is made to make the calculations easier and thus the strategy of the proof clearer.
By $S^{2}(V)$ we denote the space of all symmetric tensors of order $2$ on a vector space $V$. By $S^{2}(V)^{H}$ we denote the space of all symmetric tensors of order $2$ on a vector space $V$ which are $H$-invariant. As a consequence of assumption 3.2 we have the splitting
$S^{2}(\mathfrak{n})^{H}=S^{2}(\mathfrak{p}\oplus\mathfrak{m})^{H}=S^{2}(\mathfrak{p})^{H}\oplus S^{2}(\mathfrak{m})^{H},$
see e.g. [8].
We will determine the normal part of the tension field by making use of Theorem 2.3.
From (3.2) in [25] we have
${P_{t}}_{\lvert\mathfrak{p}}=t^{2}\operatorname{Id}_{\mathfrak{p}}+t^{4}B(t),$
${P_{t}}_{\lvert\mathfrak{m}}=A_{0}+tA_{1}+t^{2}A(t),$
where $A_{0},A_{1}$ are constant matrices and $A,B$ are smooth functions of $t$. Note that ${P_{t}}_{\lvert\mathfrak{mp}}=0$ due to the assumption (3.2). Thus we obtain
$\dot{P_{t}}_{\lvert\mathfrak{p}}=2t\operatorname{Id}_{\mathfrak{p}}+4t^{3}B(t)+t^{4}\dot{B}(t),$
$\dot{P_{t}}_{\lvert\mathfrak{m}}=A_{1}+2tA(t)+t^{2}\dot{A}(t).$
From (3.3) in [25] we have
${P_{t}}^{-1}_{\lvert\mathfrak{p}}=\frac{1}{t^{2}}\operatorname{Id}_{\mathfrak{p}}-B(t)+$...,
${P_{t}}^{-1}_{\lvert\mathfrak{m}}=A_{0}^{-1}-tA_{0}^{-1}A_{1}A_{0}^{-1}+.$$\quad.$
Thus we get
$\frac{1}{2}\operatorname{Tr}_{\mathfrak{p}}({P_{t}}^{-1}\dot{P_{t}})=\frac{1}{t}\dim\mathfrak{p}+f_{1}(t),$
$\frac{1}{2}\operatorname{Tr}_{\mathfrak{m}}({P_{t}}^{-1}\dot{P_{t}})=f_{2}(t),$
where $f_{1},f_{2}:[0,\infty)\rightarrow\mathbb{R}$ are smooth functions. Furthermore, we obtain
$\frac{1}{2}\operatorname{Tr}_{\mathfrak{p}}({P_{t}}^{-1}\dot{P_{r(t)}})=\frac{r(t)}{t^{2}}\dim\mathfrak{p}+f_{3}(t,r(t)),$
$\frac{1}{2}\operatorname{Tr}_{\mathfrak{m}}({P_{t}}^{-1}\dot{P_{r(t)}})=f_{4}(t,r(t)),$
where $f_{3},f_{4}:[0,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ are smooth functions.
We thus obtain from Theorem 2.3:
$\ddot{r}(t)+\dim\mathfrak{p}\frac{\dot{r}(t)}{t}-\dim\mathfrak{p}\frac{r(t)}{t^{2}}+f_{5}(t,r(t),\dot{r}(t))=0,$
where
$f_{5}(t,r(t),\dot{r}(t))=(f_{1}(t)+f_{2}(t))\dot{r}(t)+f_{3}(t,r(t))+f_{4}(t,r(t)).$
Clearly $f_{5}:[0,\infty)\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ is smooth.
We will deal with the general case now, i.e. we now do not longer assume that (3.2) holds. Note that for a block matrix
$X=\begin{pmatrix}A&B\\ C&D\end{pmatrix}$,
where $A$ and $D$ are square matrices. If $A$ and its Schur complement
$E:=(D-CA^{-1}B)^{-1}$
are invertible, then $X$ is invertible and the inverse of $X$ is given by
$X^{-1}=\begin{pmatrix}A^{-1}+A^{-1}BECA^{-1}&-A^{-1}BE\\ -ECA^{-1}&E\end{pmatrix}$.
(An analogous statement holds if $D$ and its Schur complement $A-BD^{-1}C$ are invertible.) We apply these results to
$P_{t}=\begin{pmatrix}{P_{t}}_{\lvert\mathfrak{p}}&{P_{t}}_{\lvert\mathfrak{mp}}\\ {P_{t}}_{\lvert\mathfrak{mp}}^{\mathrm{T}}&{P_{t}}_{\lvert\mathfrak{m}}\end{pmatrix}$.
From (3.2) and (3.3) in [25] we have
${P_{t}}_{\lvert\mathfrak{mp}}=t^{2}C_{0}+t^{3}C(t),$
${P_{t}}_{\lvert\mathfrak{mp}}^{-1}=-A_{0}^{-1}C_{0}+$...,
where $C_{0}$ is a constant matrix and $C$ is smooth. We start by determining the Schur complement $E$ of ${P_{t}}_{\lvert\mathfrak{p}}$. Plugging in the expressions for ${P_{t}}_{\lvert\mathfrak{mp}},{P_{t}}_{\lvert\mathfrak{m}},{P_{t}}_{\lvert\mathfrak{p}}$ yields
$E=({P_{t}}_{\lvert\mathfrak{m}}-{P_{t}}_{\lvert\mathfrak{mp}}^{\mathrm{T}}{P_{t}}_{\lvert\mathfrak{p}}{P_{t}}_{\lvert\mathfrak{mp}})^{-1}=A_{0}^{-1}-tA_{0}^{-1}A_{1}A_{0}^{-1}+.$$\quad.$
Analogous calculations yield
[…]
Thus straightforward computations yield
$\frac{1}{2}\operatorname{Tr}_{\mathfrak{p}}({P_{t}}^{-1}\dot{P_{t}})=\frac{1}{t}\dim\mathfrak{p}+f_{1}(t),$
$\frac{1}{2}\operatorname{Tr}_{\mathfrak{m}}({P_{t}}^{-1}\dot{P_{t}})=f_{2}(t),$
where $f_{1},f_{2}:[0,\infty)\rightarrow\mathbb{R}$ are smooth functions. Further, we obtain
$\frac{1}{2}\operatorname{Tr}_{\mathfrak{p}}({P_{t}}^{-1}\dot{P_{r(t)}})=\frac{r(t)}{t^{2}}\dim\mathfrak{p}+f_{3}(t,r(t)),$
$\frac{1}{2}\operatorname{Tr}_{\mathfrak{m}}({P_{t}}^{-1}\dot{P_{r(t)}})=f_{4}(t,r(t)),$
where $f_{3},f_{4}:[0,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ are smooth functions. Note that the functions $f_{1},f_{2},f_{3},f_{4}$ are possibly not the same as above, where we assumed (3.2). Thus, as above, we get from Theorem 2.3:
$\ddot{r}(t)+\dim\mathfrak{p}\frac{\dot{r}(t)}{t}-\dim\mathfrak{p}\frac{r(t)}{t^{2}}+f_{5}(t,r(t),\dot{r}(t))=0,$ (3.1)
where
$f_{5}(t,r(t),\dot{r}(t))=(f_{1}(t)+f_{2}(t))\dot{r}(t)+f_{3}(t,r(t))+f_{4}(t,r(t))$
and $f_{5}:[0,\infty)\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ is smooth. Hence we are left with solving the initial value problem associated to (3.1).
We rewrite (3.1) as system of first order equations:
$\dot{r}(t)=s(t),$
$\dot{s}(t)=-\dim\mathfrak{p}\frac{s(t)}{t}+\dim\mathfrak{p}\frac{r(t)}{t^{2}}-f_{5}(t,r(t),s(t)),$ (3.2)
with initial conditions $r(0)=0$ and $s(0)=v$ for some $v\in\mathbb{R}$. This system can also be rewritten as
$\dot{r}(t)=s(t),$
$\dot{s}(t)=\frac{f_{6}(t,r(t),s(t))}{t}+\dim\mathfrak{p}\frac{r(t)}{t^{2}},$ (3.3)
for $f_{6}:[0,\infty)\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ smooth.
Following [9] and [8], one can use a modified Picard iteration to show that for each $v$ there exists a unique smooth solution of (3) depending continuously on $v$. This method has been discussed in [8] for a large class of systems, in particular (3) falls into the category of systems discussed there. This machinery works since $f_{6}(t,r,s)$ is bounded and uniformly Lipschitz in the variables $r$ and $s$ under the condition that $r$ and svs-v are bounded by some sufficiently small constant, whence the claim. ∎
Remark 3.3.
  1. (1)
    Note that
    $\ddot{r}(t)+\dim\mathfrak{p}\frac{\dot{r}(t)}{t}-\dim\mathfrak{p}\frac{r(t)}{t^{2}}=0,$
    is a regular-singular ordinary differential equation.
  2. (2)
    For specific cases the initial value problem for harmonic maps of cohomogeneity one manifolds has been solved in the literature, see e.g. [7, 10, 19, 20].
  3. (3)
    The same strategy as in the present section can be used to solve the initial value problem for equivariant harmonic maps between (possibly) different cohomogeneity one manifolds.
4. Regular-singular systems of first order
In the present section we present a theory for regular-singular systems of first order. Many of the results are known, but scattered throughout the literature. The organization of the section is as follows: In the first two subsections we consider the regular-singular differential equation/system whose solutions are power functions/vector-valued generalizations of power functions. Basic phenomena are explained with the help of these examples. Afterwards, in Subsections 4.3-4.7 we study linear regular-singular system of first order. In particular, we discuss the monodromy map. Finally, in Subsections 4.8-4.10 we study non-linear regular-singular system of first order.
We refer to [13, 21, 23, 14] and the references therein.
4.1. Scalar power functions
The power function $y=c\cdot x^{a}$, with $a\in{\mathbb{C}}$, $c\in{\mathbb{C}}^{*}$, satisfies the differential equation
$\frac{dy}{dx}=\frac{a}{x}\cdot y,$ (4.1)
and is, up to a scaling factor, determined uniquely by this differential equation. Notice, that the power function $y$ is a holomorphic function in the classical sense if and only if $a\in\mathbb{Z}$. For all other choices of $a\in\mathbb{C}$, one has to slice the complex plane along one ray emanating from the origin, for example one could slice the complex plane along the negative real axis. In other words, we have to deal with a multi-valued function, which is better characterized by its graph
$\left\{(x,y)\in{\mathbb{C}}^{*}\times{\mathbb{C}}\,\lvert\,y=c\cdot x^{a}\right\}=\mbox{im}\big(z\mapsto(e^{z},c\cdot e^{az}),z\in{\mathbb{C}}\big).$
Remark 4.1.
The fact that the right hand side of (4.1) has only one pole of first order in $x=0$ is crucial for the identification of $y$ with a power function. Notice, that already slightly more complicated singularities on the right hand side of (4.1) may cause decisively more complicated essential singularities. For example, the functions
$y=c\cdot e^{-1/x},\hskip 2.84544ptx\in{\mathbb{C}}^{*}$
satisfy the differential equation
$\frac{dy}{dx}=\frac{1}{x^{2}}\cdot y,$
whose right hand side has a pole of second order at $x=0$.
4.2. Vector-valued generalizations of the scalar power functions
By the considerations of the preceding subsection it seems natural to consider the holomorphic maps $Y:U\rightarrow{\mathbb{C}}^{n}$, with open $U\subset{\mathbb{C}}^{*}$, which satisfy
$\frac{d}{ds}Y=\frac{1}{s}A_{0}\cdot Y,$ (4.2)
for a fixed $A_{0}\in\mbox{End}({\mathbb{C}}^{n})$, as vector-valued generalizations of the scalar power functions.
As in the case of (4.1), the solutions of the differential equation (4.2) can be determined explicitly. Indeed, using the projection $\widehat{\pi}$$\mathbb{C}$$\rightarrow{\mathbb{C}}^{*}$, $z\mapsto s=e^{z}$, one can pull back the differential equation to a simply-connected domain of definition. Identity (4.2) thus transforms into
[…]
and consequently one gets
$(Y\circ\widehat{\pi})(z)=\exp(z\,A_{0})\cdot(Y\circ\widehat{\pi})(0),$ (4.3)
or, less formal,
$Y(s)=\exp(\ln(s)\,A_{0})\cdot Y(1).$
Notice, that $Y:{\mathbb{C}}^{*}\rightarrow{\mathbb{C}}^{n}$ is a holomorphic function in the classical sense if and only if $A_{0}$ is diagonalizable and has only integer eigenvalues.
Otherwise, one has to deal with non-trivial monodromy. If one solves (4.2) along a circle line $c:\theta\mapsto s_{0}\,e^{i\theta}$, $0\leq\theta\leq 2\pi$, then
$\big(Y\circ c\big)(2\pi)\neq\big(Y\circ c\big)(0)$
holds in the general case. Indeed, using equation (4.3) and choosing $z_{0}\in{\mathbb{C}}$ such that $s_{0}=e^{z_{0}}$, we get
$\big(Y\circ c\big)(2\pi)=\exp(2\pi i\,A_{0})\cdot\big(Y\circ c\big)(0).$
One calls $M_{0}:=\exp(2\pi i\,A_{0})$ the monodromy generator.
4.3. Linear regular-singular system of first order in $n$ variables
One can substitute (4.2) by a more general differential equation without changing the solution theory decisively. We consider domains of the form
$\Omega_{\rho}=\left\{\begin{array}[]{ll}B(0,\rho)\subset{\mathbb{C}},&\mbox{if $0<\rho<\infty$;}\\ {\mathbb{C}},&\mbox{if $\rho=\infty$;}\end{array}\right$.
and set $\Omega_{\rho}^{\circ}=\Omega_{\rho}-\left\{0\right\}$ for $0<\rho<\infty$.
Definition 4.2.
Let $A:\Omega_{\rho}\rightarrow\mbox{End}({\mathbb{C}}^{n})$ and $h:\Omega_{\rho}\rightarrow{\mathbb{C}}^{n}$ be holomorphic maps. Then the differential equation system
$\frac{d}{ds}Y+\frac{1}{s}A(s)\cdot Y=h(s),\hskip 5.69046pts\in\Omega_{\rho}^{\circ},$ (4.4)
is called a (local) linear regular-singular system of first order in $n$ variables.
The point $s=0$ is henceforth referred to as singular point. As usual, if $h=0$ the system (4.4) is called homogeneous and inhomogeneous otherwise.
Here, too it is useful to consider the pull-back of (4.4) with respect to the covering map $\widehat{\pi}$$\mathbb{C}$$\rightarrow{\mathbb{C}}^{*}$, $z\mapsto s=e^{z}$. In other words, we deal with the differential equation system
$\frac{d}{dz}\widehat{Y}+\widehat{A}(z)\cdot\widehat{Y}=e^{z}\,\widehat{h}(z),\,z\in\widehat{\pi}^{-1}(\Omega_{\rho}^{\circ}),$ (4.5)
with $\widehat{A}(z)=A\circ\pi(z)=A(e^{z})$ and $\widehat{h}(z)=h\circ\widehat{\pi}(z)=h(e^{z})$. Since the set
$\widehat{\pi}^{-1}(\Omega_{\rho}^{\circ})=\left\{z\in{\mathbb{C}}\,\lvert\,\Re(z)<\ln(\rho)\right\}$
is a half-space, it is simply connected. Consequently, there exist global solutions $\widehat{Y}:\widehat{\pi}^{-1}(\Omega_{\rho}^{\circ})\rightarrow{\mathbb{C}}^{n}$ of (4.5). Furthermore, one can prescribe the initial value
$\widehat{Y}(z_{0})=\widehat{Y_{0}}\in{\mathbb{C}}^{n}$ (4.6)
at the arbitrarily chosen basis point $z_{0}\in\widehat{\pi}^{-1}(\Omega_{\rho}^{\circ})$. The solution of the initial value problem (4.5), (4.6) is determined uniquely.
4.4. Fundamental solution of the homogeneous system
For an explicit characterization of the solutions of (4.5),(4.6) it is useful to introduce the fundamental solution $\widehat{U}:\widehat{\pi}^{-1}(\Omega_{\rho}^{\circ})\rightarrow\mbox{End}({\mathbb{C}}^{n})$ of the associated homogeneous system
[…]
It is well-known that $U(z)\in\mbox{Gl}(n,{\mathbb{C}})\subset\mbox{End}({\mathbb{C}}^{n})$ for all $z\in\widehat{\pi}^{-1}(\Omega_{\rho}^{\circ})$.
Since $\widehat{A}(z+2\pi i)=\widehat{A}(z)$, for any solution $U$ of (4.5), (4.6) the map
$z\mapsto\widehat{U}(z+2\pi i)\cdot\widehat{U}(z_{0}+2\pi i)^{-1}$
is a solution of the initial value problem (4.5), (4.6). Consequently, we obtain
$\widehat{U}(z+2\pi i)=\widehat{U}(z)\cdot\widehat{U}(z_{0}+2\pi i).$ (4.8)
[…]
$\widehat{U}(z)^{-1}\widehat{U}(z+2\pi i)=\widehat{U}(z_{0})^{-1}\widehat{U}(z_{0}+2\pi i).$ (4.9)
In the general case the right hand side of the preceding equation does not coincide with the identity matrix. Consequently, in general we do not have a holomorphic map $U:\Omega_{\rho}^{\circ}\rightarrow\mbox{Gl}(n,{\mathbb{C}})$ with $\widehat{U}(z)=U(e^{z})$. However, the following lemma holds.
Lemma 4.3.
There exists a holomorphic map $M:\Omega_{\rho}^{\circ}\rightarrow\mbox{Gl}(n,{\mathbb{C}})$, such that the identity $\widehat{U}(z+2\pi i)\cdot\widehat{U}(z)^{-1}=M(e^{z})$ holds.
Obviously, the map $M$ is independent of the choice of the initial values in (4.7). Important is only that $\widehat{U}(z_{0})\in\mbox{Gl}(n,{\mathbb{C}})$ holds. The significance of the monodromy map $M$ is explained later on.
Proof.
Using (4.8) one obtains by a straightforward computation that $\widehat{M}(z)=\widehat{U}(z+2\pi i)\cdot\widehat{U}(z)^{-1}$ satisfies $\widehat{M}(z+2\pi i)=\widehat{M}(z)$. ∎
4.5. Solutions of the inhomogeneous equation
In the present subsection we deal with the solutions of the inhomogeneous equation (4.5). If $\widehat{h}=0$, then the map
$z\mapsto\widehat{Y}(z):=\widehat{U}(z)\cdot\widehat{Y_{0}}$
is a solution of (4.5) with the correct initial value $\widehat{Y}(z_{0})=\widehat{Y_{0}}$.
Lemma $4.$
The solution of the inhomogeneous IVP (4.5), (4.6) is given by
$\widehat{Y}(z)$ $=\widehat{U}(z)\cdot\big(\widehat{Y_{0}}+\int_{z_{0}}^{z}e^{\xi}\,\widehat{U}(\xi)^{-1}\cdot\widehat{h}(\xi)\,d\xi\big)$
$=\widehat{U}(z)\cdot\widehat{Y_{0}}+\int_{z_{0}}^{z}e^{\xi}\,\widehat{U}(z)\cdot\widehat{U}(\xi)^{-1}\cdot\widehat{h}(\xi)\,d\xi.$
Proof.
The map $V(z):=\widehat{U}(z)^{-1}\cdot\widehat{Y}(z)$ satisfies
$\frac{d}{dz}V(z)=e^{z}\,\widehat{U}(z)^{-1}\cdot\widehat{h}(z).$
Hence, the claimed identity follows from the fundamental theorem of calculus. ∎
4.6. Discussion of the monodromy map
In the present subsection we prove that the point $s=0$ is a removable singularity of the monodromy map $M$. Furthermore, we examine whether the matrices $M(s)$ and $M(s_{0})$ are conjugate for all $s\in\Omega_{\rho}$.
Using (4.9) one proves easily that
$M(e^{z})=\widehat{U}(z)\cdot M(e^{z_{0}})\cdot\widehat{U}(z)^{-1}$
and thus the matrices $M(s)$ and $M(s_{0})$ are conjugate for $s\neq 0$. Consequently, for $s\neq 0$ the characteristic polynomials of the monodromy matrices $M(s)$ coincide. The same holds for the minimal polynomials of the matrices $M(s)$ with $s\neq 0$.
Notice, that $\lim_{x\rightarrow\infty}(e^{z-x})=0$. However, the difficulty consists in the fact that in general either the limes $\lim_{x\rightarrow\infty}\widehat{U}(z-x)$ does not exist or is not an invertible matrix. We overcome these difficulties by a more subtle covering construction. In what follows we consider the differential equation
$\frac{d}{ds}\widetilde{U}+\frac{1}{s}\,A(s)\cdot\widetilde{U}=0$
[…]
If $\sigma=e^{z_{0}}$, then the map $t\mapsto\widehat{U}(z_{0}+it)$ solves the initial value problem (4.10). In other words, we get
$M(\sigma)=U(\sigma,2\pi).$
Since $A$ is holomorphic we get $A(s)=A_{0}+i\,\widetilde{A_{1}}(s)$ with $\widetilde{A_{1}}$ holomorphic. Plugging in this expression in (4.10) we get
[…]
Now we let converge $\sigma$ against 0 in the preceding initial value problem. The solution $U$ depends differentiable on $\sigma$ and it is obvious that
$U(0,t)=\exp(-it\,A_{0})$
holds. Hence, we have proven the following proposition.
Proposition 4.5.
The monodromy map $M$ is continuous and thus holomorphic continuable to $\Omega_{\rho}$. Furthermore, the identities
$M(0)=U(0,2\pi)=\exp(-2\pi i\,A_{0})$
hold.
Corollary 4.6.
The minimal polynomial of the matrices $M(s)$, $s\neq 0$, annuls $M(0)$.
Warning: The fact that $M$ is holomorphic continuable to $\Omega_{\rho}$ does not imply that $M(0)$ is conjugate to a matrix $M(\sigma)$ with $\sigma\in{\mathbb{C}}^{*}$. A well-known counterexample are the nilpotent matrices $M(\sigma)=\left(\begin{array}[]{cc}0&\sigma\\ 0&0\end{array}\right)$, which are pairwise conjugate for $\sigma\neq 0$. However, $\lim_{\sigma\rightarrow 0}M(\sigma)=0$, i.e., they converge against a matrix in a different conjugacy class.
Lemma 4.7.
The preceding counterexample also works in the context of monodromy.
Proof.
For given $\lambda\in{\mathbb{C}}$ we consider the matrix
$A(s)=A_{0}+sA_{1},\hskip 5.69046pt\mbox{with}\hskip 5.69046ptA_{0}=\left(\begin{array}[]{cc}\lambda&0\\ 0&\lambda+1\end{array}\right)\hskip 5.69046pt\mbox{and}\hskip 5.69046ptA_{1}=\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right).$
Hence,
[…]
and the differential equation system for the fundamental solution is given by
$\frac{d}{dz}\widehat{U}+A_{0}\cdot\widehat{U}+e^{z}\,A_{1}\cdot\widehat{U}=0$
or, equivalently, in terms of the map $\widehat{V}(z)=\exp(zA_{0})\cdot\widehat{U}(z)$ by
$0=\frac{d}{dz}\widehat{V}+\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)\cdot\widehat{V}$.
In other words, we get
[…]
Consequently, the fundamental solution is — up to a right multiplication with an invertible matrix — of the form
$\widehat{U}(z)=\exp(-z\,A_{0})\cdot\widehat{V}(z)=\exp(-z\,A_{0})\cdot\left(\begin{array}[]{cc}1&-z\\ 0&1\end{array}\right)=e^{-\lambda\,z}\,\left(\begin{array}[]{cc}1&-z\\ 0&e^{-z}\end{array}\right).$
Hence,
$M(e^{z})=\widehat{U}(z+2\pi i)\cdot\widehat{U}(z)^{-1}=e^{-2\pi i\,\lambda}\,\left(\begin{array}[]{cc}1&-2\pi i\,e^{z}\\ 0&1\end{array}\right),$
and thus
$M(\sigma)=e^{-2\pi i\,\lambda}\,\left(\begin{array}[]{cc}1&-2\pi i\,\sigma\\ 0&1\end{array}\right).$
Obviously, the matrix $M(0)$ is semi-simple and thus in particular not conjugate to $M(\sigma)$, $\sigma\neq 0$, since these matrices have a nilpotent part. ∎
4.7. Geometric properties of the monodromy map
The map
[…]
is fixpoint free and biholomorphic on the domain $\widehat{\pi}^{-1}(\Omega_{\rho}^{\circ})\times{\mathbb{C}}^{n}\subset{\mathbb{C}}^{n+1}$. Hence, we get a free action of $\mathbb{Z}=\left\langle\varphi\right\rangle$. Consequently, we obtain a $n$-dimensional holomorphic vector bundle
$E:=(\widehat{\pi}^{-1}(\Omega_{\rho}^{\circ})\times{\mathbb{C}}^{n})/\left\langle\varphi\right\rangle\rightarrow\Omega_{\rho}^{\circ}$
and the fundamental solution $\widehat{U}$ of (4.7) is a holomorphic trivialization of this vector bundle.
The solutions of the inhomogeneous system (4.5) are — under certain conditions — in a holomorphic $n$-dimensional affine bundle $E_{\mbox{aff}}\rightarrow\Omega_{\rho}^{\circ}$, on which the vector bundle $E\rightarrow\Omega_{\rho}^{\circ}$ acts naturally by addition
[…]
where $E\times_{\Omega_{\rho}^{\circ}}E_{\mbox{aff}}=\left\{(\widehat{Y},f)\in E_{\mbox{aff}}\,\lvert\,\pi(\widehat{Y})=\pi(f)\right\}$ denotes the fiber product of these bundles.
4.8. Non-linear regular-singular systems
Let
$0=\frac{d}{d\xi}Y+\frac{1}{\xi}\,f(\xi,Y),$ (4.11)
be a differential system defined on a punctured neighborhood $\Omega_{\rho}=B(0,\rho)-\left\{0\right\}\subset{\mathbb{C}}$ of 0 where $f:B(0,\rho)\times{\mathbb{C}}^{n}\rightarrow{\mathbb{C}}^{n}$ is a differentiable/ holomorphic function. Such a system is called a local, non-linear, regular-singular differential equation system.
Notice, that every such function $f$ possesses a expansion of the form
$f(\xi,Y)=\sum_{k=0}^{\infty}f_{k}(\xi,Y\,...\,Y),$
where we consider $f_{k}$ as holomorphic map in the $k$-linear maps, i.e., $B(0,\rho)\rightarrow L^{k}({\mathbb{C}}^{n}\times\,...\,\times{\mathbb{C}}^{n},{\mathbb{C}}^{n})$.
We call (4.11) weakly non-linear in the singularity at $z=0$ if and only if
$f_{k}(\xi,Y\,...\,Y)=\xi\,\widehat{f_{k}}(\xi,Y\,...\,Y)$
for all $k\geq 2$, where $\widehat{f_{k}}:B(0,\rho)\rightarrow L^{k}({\mathbb{C}}^{n}\times\,...\,\times{\mathbb{C}}^{n},{\mathbb{C}}^{n})$ is holomorphic.
Here, too we consider the pull back of (4.11) under $\exp$
$\widehat{\Omega_{\rho}}:=\left\{z\in{\mathbb{C}}\,\lvert\,\mbox{im}(z)<\ln(\rho)\right\}\rightarrow\Omega_{\rho},z\mapsto\xi=e^{z},$
i.e., we consider the differential equation system
$0=\frac{d}{dz}\big(Y\circ\exp\big)+f\big(e^{z},Y\circ\exp(z)\big).$
Thus we obtain a simply connected domain of definition and, furthermore, we avoid monodromy-problems and multi-valued functions.
4.8.1. Continuous continuation of the solutions to $\xi=0$
Lemma 4.8.
Let $\xi_{0}\in\Omega_{\rho}$ and let $Y$ be a solution of (4.11) along the line $t\mapsto t\xi_{0}$, $t\in(0,1]$. Furthermore, assume that the map $t\mapsto Y(t\,\xi_{0})$ is continuous continuable to $t=0$ with $\lim_{t\rightarrow 0}Y(t\,\xi_{0})=:Y_{0}$. Then,
$f(0,Y_{0})=0$
holds.
Proof.
The main theorem of calculus implies
$Y(t_{0}\,\xi_{0})-Y_{0}=\int_{0}^{t_{0}}\frac{1}{t}\,\frac{1}{\xi_{0}}f(t\,\xi_{0},Y(t\,\xi_{0}))\,dt.$
Suppose that $W_{0}:=\frac{1}{\xi_{0}}f(0,Y_{0})\neq 0$. Then, we obtain
$\lvert\frac{1}{\xi_{0}}f(t\,\xi_{0},Y(t\,\xi_{0}))-W_{0}\lvert\,\leq\,\frac{1}{2}\,\lvert W_{0}\lvert$
for all $t\in\left[0,t_{0}\right]$ under the condition, that $t_{0}\in(0,1]$ is sufficiently small. This in turn would imply
[…]
$\,\geq\frac{1}{2}\,\ln(\frac{t_{0}}{\theta})\,\lvert W_{0}\lvert$.
Consequently, we would get
$\lim_{\theta\searrow 0}\lvert Y(t_{0}\,\xi_{0})-Y(\theta\,\xi_{0})\lvert=\lim_{\theta\searrow 0}\lvert\int_{\theta}^{t_{0}}\frac{1}{t}\,\frac{1}{\xi_{0}}f(t\,\xi_{0},Y(t\,\xi_{0}))\,dt\lvert$$\infty$.
However, this contradicts the continuity of the map $\theta\mapsto\lvert Y(t_{0}\,\xi_{0})-Y(\theta\,\xi_{0})\lvert$ in $\theta=0$.
4.8.2. Normalization of continuous continuable solutions
Let $Y$ be a solution of (4.11) along the line $t\mapsto t\xi_{0}$, $t\in(0,1]$, such that the map $t\mapsto Y(t\,\xi_{0})$ is continuous continuable to $t=0$ with $\lim_{t\rightarrow 0}Y(t\,\xi_{0})=:Y_{0}$. Then the map
$\widetilde{Y}:=Y-Y_{0}$
also solves a non-linear, regular-singular system with the new right hand side
$\widetilde{f}(\xi,\widetilde{Y}):=f(\xi,\widetilde{Y}+Y_{0})$
and the corresponding expansion terms $\widetilde{f_{k}}$.
The preceding transformation has the following properties:
  1. (1)
    $\widetilde{f}(0,0)=f(0,Y_{0})=0$,
  2. (2)
    if $f$ is weakly non-linear, then $\widetilde{f}$ is also weakly non-linear and
    $\widetilde{f_{1}}(0,\widetilde{Y})=f(0,\widetilde{Y})=:A_{0}\cdot\widetilde{Y}$
    holds.
Consequently, we obtain
[…]
4.9. Smooth solutions of the non-linear regular-singular system
Proposition 4.9.
Let $Y$ be a local solution — i.e., it is defined in a punctured neighborhood of $\xi=0$ — of the regular-singular system
$0=\frac{d}{d\xi}Y+\frac{1}{\xi}\,f(\xi,Y),$ (4.12)
where $f$ is an analytic function. Then the following statements hold:
  1. (1)
    if $Y$ is continuous continuable to $\xi=0$ with $\lim_{\xi\rightarrow 0}Y(\xi)=Y_{0}$, then
    $f(0,Y_{0})=0,$
  2. (2)
    if $Y$ is in addition differentiable at the point $\xi=0$ with $\frac{d}{d\xi}Y(0)=Y_{1}$, or if $Y^{{}^{\prime}}$ is continuous continuable to $\xi=0$ with $\lim_{\xi\rightarrow 0}Y^{{}^{\prime}}(\xi)=Y_{1}$, then we get
[…]
    with $a_{0}:=\frac{\partial f}{\partial\xi}(0,Y_{0})$ and $A_{0}:=\frac{\partial f}{\partial Y}(0,Y_{0})$.
Remark 4.10.
The additional prerequisites in 2 guarantee each that the function
$\widehat{Y}(\xi):=\frac{1}{\xi}\,(Y(\xi)-Y_{0})=\int_{0}^{1}Y^{{}^{\prime}}(t\,\xi)\,dt,$
which is a priori only defined in a punctured neighborhood of $\xi=0$, is continuous continuable to $\xi=0$.
Lemma 4.11.
Let $f(0,Y_{0})=0$. Then $Y$ is a solution of (4.12) if and only if $\widehat{Y}$ is a solution of
[…]
where
[…]
with $\widehat{f_{1}}$ analytic.
Proof.
Obviously, $\frac{d}{d\xi}Y=\xi\,\frac{d}{d\xi}\widehat{Y}+\widehat{Y}$. Consequently, using $f(0,Y_{0})=0$, the identity (4.12) is equivalent to
$0=\frac{d}{d\xi}\widehat{Y}+\frac{1}{\xi}\,\big[\widehat{Y}+\int_{0}^{1}\frac{1}{\xi}\frac{\partial}{\partial t}(f(t\xi,Y_{0}+t\xi\widehat{Y}))\,dt\big].$
Then the claim follows immediately by calculating the integrand. ∎
Lemma 4.12.
If $Y$ is bounded and differentiable in a punctured neighborhood $U$ of $\xi=0$ and and if
$0=\frac{d}{d\xi}Y+\frac{1}{\xi}\,h(\xi)$
holds on $U$, where $h$ is a continuous function on a neighborhood of $\xi=0$, then
$h(0)=0.$
Proof.
Assume $h(0)\neq 0$. The continuity of $h$ in $\xi=0$ implies
$\lvert h(\xi)-h(0)\lvert\,<\,\frac{1}{2}\,\lvert h(0)\lvert$
for all $\xi\neq 0$ which are sufficiently small. Hence, we get
$Y(\xi)-Y(t\,\xi)=\int_{t}^{1}\xi\,Y^{{}^{\prime}}(\theta\,\xi)\,d\theta=-\int_{t}^{1}\frac{1}{\theta}\,h(\theta\xi)\,d\theta$
Consequently, we obtain
$\lvert Y(\xi)-Y(t\,\xi)\lvert\,$ $\,\geq\int_{t}^{1}\frac{1}{\theta}\,\lvert h(0)\lvert\,d\theta-\int_{0}^{1}\frac{1}{\theta}\,\lvert h(\theta\,\xi)-h(0)\lvert\,d\theta$
$\,\geq\lvert h(0)\lvert\,\ln(\frac{1}{t})-\frac{1}{2}\,\lvert h(0)\lvert\,\ln(\frac{1}{t})=\frac{1}{2}\,\lvert h(0)\lvert\,\ln(\frac{1}{t}).$
However, this contradicts the fact that the left hand side is bounded for all $t\in\left(0,1\right)$.
In what follows we combine the previous results to prove Proposition 4.9.
Proof.
The prerequisite in 1 guarantees that $h(\xi):=f(\xi,Y(\xi))$ is continuous in a neighborhood of $\xi=0$. Therefore 1 is a direct consequence of Lemma 4.12.
Due to Remark 4.10 and Lemma 4.11, the claim 2 is a consequence of 1. ∎
Warning: The continuity of $Y$ at $\xi=0$ in general does not imply that $\widehat{Y}$ has a continuous continuation or equivalently the existence of $Y^{{}^{\prime}}$. An easy counterexample is the differential equation
$y^{{}^{\prime}}=\frac{1}{\xi}\,y-1$
with the at $\xi=0$ continuous solution
$y(\xi)=\xi\,(1+\ln(\frac{1}{\xi})).$
However, $y$ not continuous differentiable; the first derivative $y^{\prime}(\xi)=\ln(\frac{1}{\xi})$ has an essential singularity at $\xi=0$. Furthermore, this solution has non-trivial monodromy:
$y(\xi\,e^{2\pi i\,k})=y(\xi)-2\pi i\,k\xi.$
In particular, a solution of the above differential equation which is restricted to the radial segment $t\mapsto t\,\xi$, $0\leq t\leq 1$, is not determined uniquely by the initial value $y(0)=0$.
Question 4.13.
Suppose that $Y$ is a continuous solution of (4.12). Does the existence of the derivative $Y^{\prime}(0)$ guarantee that $Y^{{}^{\prime}}$ is continuous in a neighborhood of 0?
Remark 4.14.
  1. (1)
    If (4.12) possesses a holomorphic and bounded solution which is defined on a punctured neighborhood of $\xi=0$, then this solution is holomorphic in $\xi=0$. This simply follows from Riemann’s theorem on removable singularities.
  2. (2)
    Let a solution $Y$ of (4.12) with $Y(0)=Y_{0}$ which is continuous in 0 be given. Suppose there exists a stability condition for the right hand side of (4.12) which guarantees the existence and uniqueness of such solutions $Y$ along each radial segment $t\mapsto t\,\xi$, $0<t\leq 1$. Such a stability condition is sufficient to prove that the holomorphic solution, composed from such radial solutions, does not have monodromy in the punctured neighborhood of 0.
Remark (Regularity of $f$): The function $y(t):=e^{-1/t}$, $t\in{\mathbb{C}}^{*}$, has derivative $y^{\prime}(t):=1/t^{2}\,e^{-1/t}$ and hence satisfies the differential equation
$y^{{}^{\prime}}(t)=\frac{1}{t^{2}}\,y(t).$
Since $\frac{1}{t}=-\ln y(t)$ one can write this differential equation in the following regular-singular form
$y^{\prime}(t)=\frac{1}{t}\,f(t,y)$
with $f(t,y)=y\,\ln(\frac{1}{y})=\frac{1}{2}\,y\,\ln(\frac{1}{y^{2}})$. Obviously, $f$ is continuous at $y=0$. However, $f$ is not continuous differentiable at $y=0$, this is even not the case on the real axis. Hence, in the sense of complex analysis, the point $y=0$ is an essential singularity of $f$.
Consequence: In order to develop a theory of regular-singular systems it is crucial to demand sufficiently many differentiability for $f$ or even that $f$ is analytic. It is not sufficient to demand continuity of $f$.
4.10. Solutions with given initial singular value $Y_{0}$
In what follows we enumerate several facts.
  1. (1)
    The set of the admissible singular initial values $Y_{0}$ consists of the solutions of the equation
    $f(0,Y_{0})=0.$
    See Proposition 4.9 1.
  2. (2)
    The holomorphic solutions with given $Y_{0}$ are in one-to-one correspondence to the continuous/holomorphic solutions $\widehat{Y}$ of (4.13). See Proposition 4.9 2.
  3. (3)
    The structure of the right hand side of (4.13) is much more special than that of the right hand side of (4.12). Indeed, due to Lemma 4.11 the identity (4.13) is given by
[…]
    Thus the singular term is linear in $\widehat{Y}$.
  4. (4)
[…]
  5. (5)
[…]
  6. (6)
    Each local solution $\widehat{Y}$ of (4.13) with $\widehat{Y}(0)=\widehat{Y_{0}}$ determined uniquely by the spectral condition considered in the preceding item, translates into a holomorphic solution $Y$ of (4.12) with $Y(0)=Y_{0}$ and $Y^{\prime}(0)=\widehat{Y_{0}}$.
  7. (7)
    The preceding considerations, do not exclude the existence of further solutions $Y$ of (4.12) with $\lim_{\xi\rightarrow 0}Y(\xi)=Y_{0}$ which have monodromy. It is true that the latter solutions also correspond to solutions $\widehat{Y}$ of (4.13), but we do not know about their asymptotic behavior. We only know that
    $\lim_{\xi\rightarrow 0}\xi\,\widehat{Y}(\xi)=0$
    is satisfied.

### v2

@@
 There are two types of orbits: Regular and non-regular orbits.
-Regular orbits have codimension one and their union constitutes an open dense set $M_{0}$ in $M$.
+Regular orbits (also called principal orbits) have codimension one and their union constitutes an open dense set $M_{0}$ in $M$.
 A non-regular orbit with codimension strictly greater than one is called singular orbit.
@@
 More precisely one obtains an ordinary differential equation of second order for a smooth function $r:(M/G)^{\circ}\rightarrow\mathbb{R}$ of the form
-$\ddot{r}(t)+f_{1}(t)\dot{r}(t)+f_{2}(t,r(t))=0,$ (1.3)
+$\ddot{r}(t)+h_{1}(t)\dot{r}(t)+h_{2}(t,r(t))=0,$ (1.3)
 where $h_{1}:(M/G)^{\circ}\rightarrow\mathbb{R},\,h_{2}:(M/G)^{\circ}\times\mathbb{R}\rightarrow\mathbb{R}$ are smooth maps and $(M/G)^{\circ}$ denotes the interior of $M/G$.
@@
 Assume that $N$ is the fibre of $0\in M/G$ under the natural projection map $\pi:M\rightarrow M/G$.
-For each $v$ there exists a unique smooth solution $r:(M/G)^{\circ}\cup\{0\}\rightarrow\mathbb{R}$ of (1.3) with $r(0)=0$ and $\dot{r}(0)=v$ which depends continuously on $v$.
+For each $v\in\mathbb{R}$ there exists a unique smooth solution $r:(M/G)^{\circ}\cup\{0\}\rightarrow\mathbb{R}$ of (1.3) with $r(0)=0$ and $\dot{r}(0)=v$ which depends continuously on $v$.
 A precise formulation of this result is contained in Section 4.
@@
 [8, 11, 21, 22].
+In 5 we study the initial value problem for equivariant biharmonic maps between cohomogeneity one manifolds: In the last few decades the study of higher order energy functionals which extend the energy functional of a map became an active reasearch area within geometric analysis, see e.g.
+the recent manuscript [7] and the references therein.
+The presentation of this paragraph closely follows the one in [7].
+The bienergy functional, which is defined by
+$E_{2}(\phi):=\frac{1}{2}\int_{M}|\tau(\phi)|^{2}\text{ }dv,$ (1.4)
+is the most prominent generalization of the energy functional.
+Critical points of the bienergy functional $E_{2}$ are called biharmonic maps and are characterized by the vanishing of the bitension field $\tau_{2}$.
+The bitension field $\tau_{2}$ is given by
+$\tau_{2}(\phi):=\bar{\Delta}\tau(\phi)+\textrm{Tr}_{g}R^{N}(\tau(\phi),d\phi)d\phi,$ (1.5)
+where $\bar{\Delta}$ denotes the rough Lapacian acting on sections of the pullback bundle $\phi^{\ast}TN$.
+Since harmonic maps are trivial examples of biharmonic maps, we will henceforth we will restrict our considerations to non-harmonic biharmonic maps, so called proper biharmonic maps.
+In the following, we always assume that we are studying proper biharmonic maps without always writing the adjective ‘proper’.
+For equivariant maps $\varphi$ between cohomogeneity one manifolds equation the equation $\tau_{2}(\phi)=0$, where $\tau_{2}$ is given as in (1.5), reduces to an ordinary differential equation of fourth order for a smooth function $r:(M/G)^{\circ}\rightarrow\mathbb{R}$ of the form
+$r^{(4)}(t)+h_{1}(t)r^{(3)}(t)+h_{2}(t,r(t),\dot{r}(t),\ddot{r}(t))=0,$ (1.6)
+where $h_{1}:(M/G)^{\circ}\rightarrow\mathbb{R},\,h_{2}:(M/G)^{\circ}\times\mathbb{R}^{3}\rightarrow\mathbb{R}$ are smooth maps.
+We solve the initial value problem for equivariant biharmonic maps between cohomogeneity one manifolds under the additional assumption that the so-called metric endomorphism $P_{t}$, see Section 2, is diagonal.
+The metric endomorphism describes the metric of the cohomogeneity one manifold with respect to a fixed biinvariant metric on the Lie algebra $\mathfrak{g}$ of $G$.
+Theorem 1.2.
+Let a cohomogeneity one manifold $M$ with a singular orbit $N$ be given such that the metric endomorphism $P_{t}$ is diagonal.
+Assume that $N$ is the fibre of $0\in M/G$ under the natural projection map $\pi:M\rightarrow M/G$.
+For each $v=(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}$ there exists a unique smooth solution $r:(M/G)^{\circ}\cup\{0\}\rightarrow\mathbb{R}$ of (1.6) with $r(0)=0$, $\dot{r}(0)=v_{1}$, $\ddot{r}(0)=v_{2}$ and $\dddot{r}(0)=v_{3}$ which depends continuously on $v$.
+A precise formulation of this result is contained in Section 5.
 In Section 6 we study regular-singular systems of first order.
@@
 The initial value problem for harmonic maps of cohomogeneity one manifolds is set up and solved in Section 4.
-In Section 4 we study regular-singular systems of first order.
+The initial value problem for biharmonic maps of (specific) cohomogeneity one manifolds is set up and solved in Section 5.
+Finally, in Section 6, we study regular-singular systems of first order.
 2. Preliminaries
@@
 where $A_{0}$, $A_{1}$ and $C_{0}$ are constant matrices and $A,B,C$ are smooth matrix valued functions of $t$, see [27].
-2.3. Harmonic maps between cohomogeneity one manifolds
-We study equivariant harmonic maps between cohomogeneity one manifolds.
+2.3. Harmonic self-maps between cohomogeneity one manifolds
+We study equivariant harmonic self-maps between cohomogeneity one manifolds.
 Again we borrow notation and inspiration for presentation from [19].
@@
 $\psi:M\rightarrow M,\,g\cdot\gamma(t)\mapsto g\cdot\gamma(r(t)).$ (2.2)
-Here $r:M/G\rightarrow\mathbb{R}$ is a smooth function which in addition satisfies boundary conditions in cases (3) and (4). Namely, $r$ satisfies
+Here $r:(M/G)^{\circ}\rightarrow\mathbb{R}$ is a smooth function which in addition satisfies boundary conditions in cases (3) and (4). Namely, $r$ satisfies
   1. (i)
@@
 Here, $E_{1},\ldots,E_{n}\in\mathfrak{n}$ and $F_{1},\ldots,F_{n}\in\mathfrak{n}$ are such that $E^{\ast}_{1|\gamma(t)}$,…, $E^{\ast}_{n|\gamma(t)}$ form an orthonormal basis of $T_{\gamma(t)}(G\cdot\gamma(t))$ and $F^{\ast}_{1|\gamma(r(t))},\ldots,F^{\ast}_{n|\gamma(r(t))}$ form an orthonormal basis of $T_{\gamma(r(t))}(G\cdot\gamma(r(t)))$.
+3. Harmonic maps between cohomogeneity one manifolds
+In this section we determine the tension field of equivariant harmonic maps between cohomogeneity one manifolds.
+In contrast to [19] we do not assume that we necessarily consider self-maps.
+However, the considerations of this section clearly follow along the lines of [19], from where we also borrow the notation.
+Let $G$ and $K$ act with cohomogeneity one on $M$ and $N$, respectively. Further, let
+$A:G\rightarrow K$
+be a group homomorphism. Fix an unit speed normal geodesic $\gamma$ on $M$ and an unit speed normal geodesic $\hat{\gamma}$ on $N$.
+Below we let $\psi$ be the equivariant map given by
+$\psi(g\cdot\gamma(t))=A(g)\cdot\hat{\gamma}(r(t)),$
+where $r:(M/G)^{\circ}\rightarrow\mathbb{R}$ is a smooth function which in addition satisfies boundary conditions in cases (3) and (4), see (i) and (ii) in Subsection 2.3.
+We will now compute the tension field of $\psi$ and we start by computing the normal component thereof. The derivative of $\psi$ in normal direction is given by
+$d\psi_{|\gamma(t)}\cdot\dot{\gamma}(t)=\frac{d}{dt}(\psi\circ\gamma(t))=\dot{r}(t)\dot{\hat{\gamma}}(r(t)).$
+Next we compute the derivative of $\psi$ in tangential directions. For this purpose we use that the map
+$\mathfrak{g}/\mathfrak{h}\to T_{p}(G\cdot p),X\mapsto X^{\ast}_{|p}$
+is a vector space isomorphism for regular points $p$. We then obtain the following identity for the derivative of $\psi$ in tangential directions
+$d\psi_{|\gamma(t)}\cdot X^{\ast}_{|\gamma(t)}=\frac{d}{ds}\psi(\exp sX\cdot\gamma(t))_{|s=0}=\frac{d}{ds}\bigl(\exp s\,dA_{e}X\cdot\hat{\gamma}(r(t))\bigr)_{|s=0}=(dA_{e}X)^{\ast}_{|\hat{\gamma}(r(t))}.$
+Here $e$ denotes the identity element of $G$.
+The tension field $\tau$ of $\psi$ is defined by (compare the introduction)
+$\tau_{|p}=\sum_{\mu=0}^{n}\nabla d\psi_{|p}(e_{\mu},e_{\mu})=\sum_{\mu=0}^{n}\Bigl(\nabla_{e_{\mu}}(d\psi\cdot e_{\mu})-d\psi\cdot\nabla_{e_{\mu}}e_{\mu}\Bigr)_{|p},$ (3.1)
+where the vectors $e_{0},\ldots,e_{n}$ form any orthonormal basis of $T_{p}M$ and can be extended arbitrarily to vector fields on a neighborhood of $p$.
+The equivariance of the tension field implies that it suffices to evaluate the expression along $\gamma(t)$.
+By $T$ we denote the unit normal field to the principal orbits of $M$ given by $T_{|g\cdot\gamma(t)}=g\cdot\dot{\gamma}(t)$.
+Moreover, we set $e_{0}=\dot{\gamma}(t)$ and choose $E_{1},\ldots,E_{n}\in\mathfrak{g}$ such that $e_{1}=E^{\ast}_{1|\gamma(t)},\ldots,e_{n}=E^{\ast}_{n|\gamma(t)}$ form an orthonormal basis of $T_{\gamma(t)}(G\cdot\gamma(t))$.
+We get
+$\nabla_{e_{0}}(d\psi\cdot e_{0})_{|\gamma(t)}=\frac{\nabla}{dt}\bigl(\dot{r}(t)\dot{\hat{\gamma}}(r(t))\bigr)=\ddot{r}(t)\dot{\hat{\gamma}}(r(t))$
+and $(d\psi\cdot\nabla_{e_{0}}e_{0})_{|\gamma(t)}=d\psi_{|\gamma(t)}\cdot\frac{\nabla}{dt}\dot{\hat{\gamma}}(t)=0$. Hence, the tension field definition (3.1) becomes
+$\tau_{|\gamma(t)}=\ddot{r}(t)\dot{\hat{\gamma}}(r(t))+\sum_{\mu=1}^{n}\Bigl(\nabla_{(dA_{e}E_{\mu})^{\ast}}(dA_{e}E_{\mu})^{\ast}_{|\hat{\gamma}(r(t))}-d\psi_{|\gamma(t)}\cdot\nabla_{E^{\ast}_{\mu}}E^{\ast}_{\mu|\gamma(t)}\Bigr).$ (3.2)
+In order to present the normal component of the tension field in a computational convenient way, we need to introduce some notation first: Following [19], we let
+$\Pi_{t}^{r(t)}:T_{\hat{\gamma}(t)}(K\cdot\hat{\gamma}(t))\to T_{\hat{\gamma}(r(t))}(K\cdot\hat{\gamma}(r(t)))$
+denote the parallel transport along the normal geodesic $\hat{\gamma}$.
+Further, we denote by $J_{t}^{r(t)}$ the endomorphism of $T_{\hat{\gamma}(t)}(K\cdot\hat{\gamma}(t))$ given by composing the action field homomorphism $X^{\ast}_{|\hat{\gamma}(t)}\mapsto X^{\ast}_{|\hat{\gamma}(r(t))}$ with $(\Pi_{t}^{r(t)})^{-1}=\Pi_{r(t)}^{t}$.
+By $\hat{T}$ we denote the unit normal field to the principal orbits of $N$ given by $\hat{T}_{|k\cdot\hat{\gamma}(t)}=k\cdot\dot{\hat{\gamma}}(t)$, for $k\in K$.
+Moreover, we set $f_{0}=\dot{\hat{\gamma}}(t)$ and choose $F_{1},\ldots,F_{m}\in\mathfrak{k}$ such that $f_{1}=F^{\ast}_{1|\hat{\gamma}(t)},\ldots,f_{m}=F^{\ast}_{m|\hat{\gamma}(t)}$ form an orthonormal basis of $T_{\hat{\gamma}(t)}(K\cdot\hat{\gamma}(t))$.
+With this preparation at hand we can now provide the normal component of the tension field.
+Theorem 3.1.
+The normal component $\tau^{\mathrm{nor}}_{|\gamma(t)}=\langle\tau_{|\gamma(t)},\dot{\hat{\gamma}}(r(t))\rangle$ of the tension field is given by
+$\tau^{\mathrm{nor}}_{|\gamma(t)}=\ddot{r}(t)-\dot{r}(t)\operatorname{Tr}S_{|\gamma(t)}+\sum_{\mu=1}^{n}\sum_{i,j=1}^{m}c_{\mu,i}c_{\mu,j}\langle(J_{t}^{r(t)})^{\ast}(\Pi_{t}^{r(t)})^{-1}\hat{S}_{|\hat{\gamma}(r(t))}\Pi_{t}^{r(t)}J_{t}^{r(t)}f_{i},f_{j}\rangle_{\lvert\hat{\gamma}(t)},$
+where
+$c_{\mu,i}:=c_{\mu,i}(t):=\langle(dA_{e}E_{\mu})^{\ast},f_{i}\rangle_{\lvert\hat{\gamma}(t)}.$
+Moreover, $S_{|\gamma(t)}$ denotes the shape operator of the orbit $G\cdot\gamma(t)$ at $\gamma(t)$, $\hat{S}_{|\hat{\gamma}(t)}$ denotes the shape operator of the orbit $K\cdot\hat{\gamma}(t)$ at $\hat{\gamma}(t)$ and $(J_{t}^{r(t)})^{\ast}$ denotes the adjoint endomorphism of $J_{t}^{r(t)}$.
+Proof.
+We evaluate the tension field formula (3.2) and make use of the identity
+$\langle\nabla_{X^{\ast}}Y^{\ast},T\rangle=\langle X^{\ast},S\cdot Y^{\ast}\rangle,$
+which holds by the definition of the shape operator. Consequently,
+$\langle\sum_{\mu=1}^{n}\nabla_{E^{\ast}_{\mu}}E^{\ast}_{\mu},T\rangle_{|\gamma(t)}=\sum_{\mu=1}^{n}\langle e_{\mu},S_{|\gamma(t)}e_{\mu}\rangle=\operatorname{Tr}S_{|\gamma(t)}.$
+Furthermore, we obtain
+$\langle\sum_{\mu=1}^{n}\nabla_{(dA_{e}E_{\mu})^{\ast}}(dA_{e}E_{\mu})^{\ast},\hat{T}\rangle_{|\hat{\gamma}(r(t))}=\sum_{\mu=1}^{n}\langle(dA_{e}E_{\mu})^{\ast}_{|\hat{\gamma}(r(t))},\hat{S}_{|\hat{\gamma}(r(t))}(dA_{e}E_{\mu})^{\ast}_{|\hat{\gamma}(r(t))}\rangle\\ =\sum_{\mu=1}^{n}\langle J_{t}^{r(t)}(dA_{e}E_{\mu})^{\ast}_{|\hat{\gamma}(t)},(\Pi_{t}^{r(t)})^{-1}\hat{S}_{|\hat{\gamma}(r(t))}\Pi_{t}^{r(t)}J_{t}^{r(t)}(dA_{e}E_{\mu})^{\ast}_{|\hat{\gamma}(t)}\rangle.$
+Plugging
+$(dA_{e}E_{\mu})^{\ast}_{|\hat{\gamma}(t)}=\sum_{i=1}^{m}c_{\mu,i}{F_{i}^{\ast}}_{|\hat{\gamma}(t)}$
+into the previous equation yields the claim. ∎
+Remark 3.2.
+Theorem 3.1 generalizes Theorem A in [19] and recovers Theorem A when $M=N$, $G=K$ and $A=\operatorname{Id}$.
+As in [19] we provide a second formula for the normal component of the tension field, namely a formula in terms of data of the acting groups $G$ and $K$.
+Let $\hat{Q}$ be a fixed biinvariant metric on $K$.
+Denote the orthonormal complement of the Lie algebra of the principal isotropy group in $\mathfrak{k}$ by $\hat{\mathfrak{n}}$.
+Define the metric endomorphisms $\hat{P}_{t}:\hat{\mathfrak{n}}\to\hat{\mathfrak{n}}$ by
+$\hat{Q}(X,\hat{P}_{t}\cdot Y)=\langle X^{\ast},Y^{\ast}\rangle_{|\hat{\gamma}(t)}.$
+We have
+$\langle X^{\ast},\hat{S}\cdot X^{\ast}\rangle_{|\hat{\gamma}(r(t))}=-\langle X^{\ast},\nabla_{\hat{T}}X^{\ast}\rangle_{|\hat{\gamma}(r(t))}=-\tfrac{1}{2\dot{r}(t)}\tfrac{d}{dt}\langle X^{\ast},X^{\ast}\rangle_{|\hat{\gamma}(r(t))}\\ =-\tfrac{1}{2}\hat{Q}(X,(\dot{\hat{P}})_{r(t)}X)=-\tfrac{1}{2}\langle X^{\ast},(\hat{P}_{t}^{-1}(\dot{\hat{P}})_{r(t)}X)^{\ast}\rangle_{|\hat{\gamma}(t)}$
+and hence the following statement holds.
+Theorem 3.3.
+The normal component of the tension field is given by
+$\tau^{\mathrm{nor}}_{|\gamma(t)}=\ddot{r}(t)+\tfrac{1}{2}\dot{r}(t)\operatorname{Tr}P_{t}^{-1}\dot{P}_{t}-\tfrac{1}{2}\sum_{\mu=1}^{n}\sum_{i,j=1}^{m}c_{\mu,i}c_{\mu,j}P_{i,j},$
+where
+$c_{\mu,i}:=c_{\mu,i}(t):=\langle(dA_{e}E_{\mu})^{\ast},f_{i}\rangle_{\lvert\hat{\gamma}(t)}.$
+and $P_{i,j}$ is the $(i,j)$-entry of the representing matrix of $\hat{P}_{t}^{-1}(\dot{\hat{P}})_{r(t)}$.
+Remark 3.4.
+Theorem 3.3 generalizes Theorem 3.4 in [19] and recovers Theorem 3.4 when $M=N$, $G=K$ and $A=\operatorname{Id}$.
+The computation of the tangential component of the tension field is analogous to the computation in [19][Theorem 3.6] and is therefore omitted. We obtain:
+Theorem 3.5.
+The tangential component of the tension field is given by
+$\tau^{\tan}_{|\gamma(t)}=\bigl(\hat{P}_{r(t)}^{-1}\sum_{\mu=1}^{n}[dA_{e}E_{\mu},\hat{P}_{r(t)}dA_{e}E_{\mu}]\bigr)^{\ast}_{|\hat{\gamma}(r(t))}$
+where $E_{1},\ldots,E_{m}\in\mathfrak{n}$ are such that $E^{\ast}_{1|\gamma(t)},\ldots,E^{\ast}_{n|\gamma(t)}$ form an orthonormal basis of $T_{\gamma(t)}(G\cdot{\gamma}(t))$.
 4. The initial value problem for harmonic maps
@@
 $0=\ddot{r}(t)-\dot{r}(t)\operatorname{Tr}S_{|\gamma(t)}+\operatorname{Tr}\,(J_{t}^{r(t)})^{\ast}(\Pi_{t}^{r(t)})^{-1}S_{|\gamma(r(t))}\Pi_{t}^{r(t)}J_{t}^{r(t)}$
-has a smooth solution $r:(M/G)^{\circ}\cup\{0\}\to\mathbb{R}$ with $\lim_{t\to 0}r(t)=0$ and $\lim_{t\to 0}\dot{r}(t)=v$, which depends continously on $v$.
+has a smooth solution $r:(M/G)^{\circ}\cup\{0\}\to\mathbb{R}$ with $\lim_{t\to 0}r(t)=0$ and $\lim_{t\to 0}\dot{r}(t)=v$, which depends continuously on $v$.
 Before proving Theorem 4.1 we recall the following: We consider metrics of the form (2.1) and have set
@@
 Following [10] and [9], one can use a modified Picard iteration to show that for each $v$ there exists a unique smooth solution of (4) depending continuously on $v$.
-This method has been discussed in [8] for a large class of systems, in particular (3) falls into the category of systems discussed there.
+This method has been discussed in [9] for a large class of systems, in particular (5.1) falls into the category of systems discussed there.
 This machinery works since $f_{6}(t,r,s)$ is bounded and uniformly Lipschitz in the variables $r$ and $s$ under the condition that $r$ and svs-v are bounded by some sufficiently small constant, whence the claim.
@@
   3. (3)
+    The results of Theorem 5.2 can be applied to solve the initial value problem for harmonic maps between so-called models, see [16], in particular equation (2.7), for details.
+  4. (4)
     The same strategy as in the present section can be used to solve the initial value problem for equivariant harmonic maps between (possibly) different cohomogeneity one manifolds.
+5. The initial value problem for biharmonic maps
+In this section we set up and solve the initial value problem for equivariant biharmonic maps of cohomogeneity one manifolds $M$ under the assumption that the metric endomorphism $P_{t}$ is diagonal.
+We will assume that a singular orbit $N$ exists, otherwise the initial value problem is a regular ODE.
+Thus $M/G$ is either isometric to $[0,1]$ or $0,\infty)$.
+Below we set up the initial value problem at $t=0$.
+In [4] biharmonic self-maps of cohomogeneity one manifolds has been studied. For $P_{t}$ diagonal the following result has been obtained:
+Theorem 5.1 ([4]).
+For a cohomogeneity one manifold with diagonal metric, the biharmonic equation of a $(k,r)$-map reduces to the following boundary value problem:
+$\begin{cases}F=\ddot{r}(t)+\tfrac{1}{2}\dot{r}(t)\operatorname{trace}P_{t}^{-1}\dot{P}_{t}-\tfrac{1}{2}\operatorname{trace}P_{t}^{-1}\dot{P}_{r(t)}\\[4.30554pt] \ddot{F}+\dot{F}\tfrac{1}{2}\operatorname{trace}P_{t}^{-1}\dot{P}_{t}-F\tfrac{1}{2}\operatorname{trace}P_{t}^{-1}\ddot{P}_{r(t)}=0\end{cases}$
+with
+$\lim_{t\rightarrow 0}r(t)=0\quad\text{and}\quad\lim_{t\rightarrow L}r(t)=kL.$
+Clearly, we obtain the same ODE if the porbit space $M/G$ is isometric to $0,\infty)$. In this section we solve the initial value problem for biharmonic maps at singular orbits:
+Theorem 5.2.
+Let $(M,g)$ be a cohomogeneity one manifold with singular orbit, i.e. $M/G$ is isometric to a closed interval $[0,L]$ or a half line $0,\infty)$. For any $v\in\mathbb{R}^{3}$ the differential equation
+$\begin{cases}F=\ddot{r}(t)+\tfrac{1}{2}\dot{r}(t)\operatorname{trace}P_{t}^{-1}\dot{P}_{t}-\tfrac{1}{2}\operatorname{trace}P_{t}^{-1}\dot{P}_{r(t)}\\[4.30554pt] \ddot{F}+\dot{F}\tfrac{1}{2}\operatorname{trace}P_{t}^{-1}\dot{P}_{t}-F\tfrac{1}{2}\operatorname{trace}P_{t}^{-1}\ddot{P}_{r(t)}=0\end{cases}$
+has a smooth solution $r:(M/G)^{\circ}\cup\{0\}\to\mathbb{R}$ with $\lim_{t\to 0}r(t)=0$, $\lim_{t\to 0}\dot{r}(t)=v_{1}$, $\lim_{t\to 0}\ddot{r}(t)=v_{2}$ and $\lim_{t\to 0}\dddot{r}(t)=v_{3}$ which depends continuously on $v$.
+Proof.
+The strategy of the proof is the same as in Theorem 4.1. Recall that from (3.2) in [27] we have
+${P_{t}}_{\lvert\mathfrak{p}}=t^{2}\operatorname{Id}_{\mathfrak{p}}+t^{4}B(t),$
+${P_{t}}_{\lvert\mathfrak{m}}=A_{0}+tA_{1}+t^{2}A(t),$
+where $A_{0},A_{1}$ are constant matrices and $A,B$ are smooth functions of $t$. Straightforward computations give
+$\tfrac{1}{2}\operatorname{trace}P_{t}^{-1}\ddot{P}_{r(t)}=\frac{\dim\mathfrak{p}}{t^{2}}+f_{1}(t,r(t)),$
+where $f_{1}:[0,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ is smooth.
+Let the functions $s,u:[0,\infty)\rightarrow\mathbb{R}$ be defined by $s(t):=\dot{r}(t)$ and $u(t):=\dot{F}(t)$.
+Straightforward computations yield the following first oder system of differential equations:
+$\dot{r}(t)=s(t),$ (5.1)
+$\dot{s}(t)=F(t)-\frac{\dim\mathfrak{p}}{t}s(t)+s(t)f_{2}(t)+\dim\mathfrak{p}\frac{r(t)}{t^{2}}+f_{2}(t,r(t)),$
+$\dot{F}(t)=u(t),$
+$\dot{u}(t)=-\dim\mathfrak{p}\frac{u(t)}{t}+u(t)f_{3}(t)+\dim\mathfrak{p}\frac{F(t)}{t^{2}}+F(t)f_{4}(t,r(t)),$
+where $f_{2},f_{4}:[0,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ and $f_{3}:[0,\infty)\rightarrow\mathbb{R}$ are smooth.
+Following again [10] and [9], one can use a modified Picard iteration to show that for each $v$ there exists a unique smooth solution of (5.1) depending continuously on $v$. ∎
+Remark 5.3.
+The results of Theorem 5.2 can be applied to solve the initial value problem for biharmonic maps between so-called models, see [16], in particular Remark 2.4, for details.
 6. Regular-singular systems of first order
@@
 Basic phenomena are explained with the help of these examples.
-Afterwards, in Subsections 4.3-4.7 we study linear regular-singular system of first order.
+Afterwards, in Subsections 6.3-6.7 we study linear regular-singular system of first order.
 In particular, we discuss the monodromy map.
-Finally, in Subsections 4.8-4.10 we study non-linear regular-singular system of first order.
+Finally, in Subsections 6.8-6.10 we study non-linear regular-singular system of first order.
 We refer to [14, 23, 25, 15] and the references therein.
@@
 is a solution of (6.5) with the correct initial value $\widehat{Y}(z_{0})=\widehat{Y_{0}}$.
-Lemma $4.$
+Lemma 6.4.
 The solution of the inhomogeneous IVP (6.5), (6.6) is given by
@@
 In particular, a solution of the above differential equation which is restricted to the radial segment $t\mapsto t\,\xi$, $0\leq t\leq 1$, is not determined uniquely by the initial value $y(0)=0$.
-Question 4.13.
+Question 6.13.
 Suppose that $Y$ is a continuous solution of (6.12). Does the existence of the derivative $Y^{\prime}(0)$ guarantee that $Y^{{}^{\prime}}$ is continuous in a neighborhood of 0?

### v3

@@
 Organization: In Section 2 we provide preliminaries.
+We provide the tension field for equavariant smooth maps between cohomogeneity one manifolds in Section 3.
 The initial value problem for harmonic maps of cohomogeneity one manifolds is set up and solved in Section 4.
@@
 where $E_{1},\ldots,E_{m}\in\mathfrak{n}$ are such that $E^{\ast}_{1|\gamma(t)},\ldots,E^{\ast}_{n|\gamma(t)}$ form an orthonormal basis of $T_{\gamma(t)}(G\cdot{\gamma}(t))$.
+We apply the preceding results to study conformal changes of the domain metric.
+Lemma 3.6.
+Let $(M,g)$ be a cohomogeneity one manifold. Further, let $\psi:(M,\exp(2\alpha(t))g)\rightarrow(M,g)$ be the equivariant map given by
+$\psi(g\cdot\gamma(t))=g\cdot{\gamma}(r(t)),$
+where $r:(M/G)^{\circ}\rightarrow\mathbb{R}$ which in addition satisfies boundary conditions in cases (3) and (4). Then we have
+$\tau^{\mathrm{nor}}_{|\gamma(t)}=\ddot{r}(t)+\tfrac{1}{2}\dot{r}(t)\operatorname{Tr}P_{t}^{-1}\dot{P}_{t}+\dim(\mathfrak{n})\dot{\alpha}(t)\dot{r}(t)-\tfrac{1}{2}\operatorname{Tr}P_{t}^{-1}(\dot{P})_{r(t)}.$
+Proof.
+This follows from Theorem 3.3. ∎
 4. The initial value problem for harmonic maps
@@
 [13, 28, 27, 20].
+We will reduce the initial value problem to an initial value problem for which a solution theory exists. The following theorem will be applied:
+Theorem 4.2 ([11]).
+Consider the singular initial value problem
+$\dot{y}=\frac{1}{t}M_{-1}(y)+M(t,y),\qquad y(0)=y_{0},$ (4.1)
+where $y$ takes values in $\mathbb{R}^{k}$, $M_{-1}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{k}$ is a smooth function of $y$ in a neighborhood of $y_{0}$ and $M:\mathbb{R}\times\mathbb{R}^{k}\rightarrow\mathbb{R}^{k}$ is smooth in $t,y$ in a neighborhood of $(0,y_{0})$.
+Assume that
+  1. (1)
+    $M_{-1}(y_{0})=0$;
+  2. (2)
+    $h\text{Id}-d_{y_{0}}M_{-1}$ is invertible for all $h\in\mathbb{N}$, $h\geq 1$.
+Then there exists a unique solution $y(t)$ of (4.1). Furthermore $y$ depends continuously on $y_{0}$ satisfying (i) and (ii).
 We are now ready to prove Theorem 4.1.
@@
 and $f_{5}:[0,\infty)\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ is smooth. Hence we are left with solving the initial value problem associated to (4.2).
-We rewrite (4.1) as system of first order equations:
-$\dot{r}(t)=s(t),$
-$\dot{s}(t)=-\dim\mathfrak{p}\frac{s(t)}{t}+\dim\mathfrak{p}\frac{r(t)}{t^{2}}-f_{5}(t,r(t),s(t)),$ (4.2)
-with initial conditions $r(0)=0$ and $s(0)=v$ for some $v\in\mathbb{R}$. This system can also be rewritten as
-$\dot{r}(t)=s(t),$
-$\dot{s}(t)=\frac{f_{6}(t,r(t),s(t))}{t}+\dim\mathfrak{p}\frac{r(t)}{t^{2}},$ (4.3)
-for $f_{6}:[0,\infty)\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ smooth.
-Following [10] and [9], one can use a modified Picard iteration to show that for each $v$ there exists a unique smooth solution of (4) depending continuously on $v$.
-This method has been discussed in [9] for a large class of systems, in particular (5.1) falls into the category of systems discussed there.
-This machinery works since $f_{6}(t,r,s)$ is bounded and uniformly Lipschitz in the variables $r$ and $s$ under the condition that $r$ and svs-v are bounded by some sufficiently small constant, whence the claim.
-∎
-Remark 4.3.
-  1. (1)
 Note that
 $\ddot{r}(t)+\dim\mathfrak{p}\frac{\dot{r}(t)}{t}-\dim\mathfrak{p}\frac{r(t)}{t^{2}}=0,$
-    is a regular-singular ordinary differential equation.
+is a regular-singular ordinary differential equation whose solutions are given by
+$r(t)=c_{0}t+c_{1}t^{-\dim\mathfrak{p}},$
+where $c_{0},c_{1}\in\mathbb{R}$. Clearly, the solution $r$ extends smoothly to $t=0$ if and only if $c_{1}=0$.
+We thus make the Ansatz
+$r(t)=c_{0}ta(t),$ (4.3)
+where $a:(0,\infty)\rightarrow\mathbb{R}$ is smooth. We plug this Ansatz into (4.2) and hence we obtain
+$a^{\prime\prime}(t)+\frac{p+2}{t}a^{\prime}(t)=f(t,a(t),a^{\prime}(t)),$
+where $f:[0,\infty)\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ is a smooth map. We let $u(t)=a^{\prime}(t)$ and thus obtain the following first order system
+$\frac{d}{dt}\begin{pmatrix}a(t)\\ u(t)\end{pmatrix}=\frac{1}{t}\begin{pmatrix}0&0\\ 0&-(p+2)\end{pmatrix}\begin{pmatrix}a(t)\\ u(t)\end{pmatrix}+\begin{pmatrix}u(t)\\ f(t,a(t),u(t))\end{pmatrix}$.
+(4.4)
+We observe that
+$h\operatorname{Id}_{2}-\begin{pmatrix}0&0\\ 0&-(p+2)\end{pmatrix}$
+is invertible for each $h\in\mathbb{N}=\{1,2,3\dots\}$ and let $a(0)=c_{1}$, $u(0)=0$. Hence Theorem 4.2 yields that there exists a unique solution
+$y(t)=\begin{pmatrix}a(t)\\ u(t)\end{pmatrix}$
+of 4.4 with
+$y(0)=\begin{pmatrix}c_{1}\\ 0\end{pmatrix}$.
+Furthermore $y$ depends continuously on $c_{1}$. Combining with (4.3) yields the claim. ∎
+Remark 4.4.
   1. (1)
@@
 $\lim_{t\rightarrow 0}r(t)=0\quad\text{and}\quad\lim_{t\rightarrow L}r(t)=kL.$
-Clearly, we obtain the same ODE if the porbit space $M/G$ is isometric to $0,\infty)$. In this section we solve the initial value problem for biharmonic maps at singular orbits:
+Clearly, we obtain the same ODE if the orbit space $M/G$ is isometric to $0,\infty)$. In this section we solve the initial value problem for biharmonic maps at singular orbits:
 Theorem 5.2.
@@
 where $f_{2},f_{4}:[0,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ and $f_{3}:[0,\infty)\rightarrow\mathbb{R}$ are smooth.
-Following again [10] and [9], one can use a modified Picard iteration to show that for each $v$ there exists a unique smooth solution of (5.1) depending continuously on $v$. ∎
+We proceed as in the proof of Theorem 4.1 to establish the claim. ∎
 Remark 5.3.
@@
 Finally, in Subsections 6.8-6.10 we study non-linear regular-singular system of first order.
-We refer to [14, 23, 25, 15] and the references therein.
+We refer to [10, 11, 15, 16, 24, 26, 9] and the references therein.
 6.1. Scalar power functions